# Reviving Thorup’s Shortcut Conjecture

### v1

@@ 4.2 Exact Shortest Paths via $(1+\epsilon)$-Approximate Distance Shortcuts
We first present an algorithm for computing $(1+o(1/\log n))$-approximate SSSP in an $O(m)$-edge, $O(n)$-vertex directed graph with 11 call to $\mathcal{A}$} and additional $\widetilde{O}\left(m+m^{\prime}\right)$ time and $\widetilde{O}\left(h\right)$ depth. We then show how to solve exact SSSP of an $m$-edge $n$-vertex graph using $\widetilde{O}\left(1\right)$ calls to the approximation algorithm with additional $\widetilde{O}\left(m\right)$ work and $\widetilde{O}\left(h\right)$ depth. Combining the two reductions finishes the proof.
@@
We can compute an $h$-limited $(1+o(1/\log n))$-approximate shortest path distances $\{\tilde{d}^{G^{\prime}}(u)\}_{u\in V}$ in $\widetilde{O}\left(m+m^{\prime}\right)$ work and $\widetilde{O}\left(h\right)$ depth (by setting the parameter $p=m/k$ in their paper). We use $\{\tilde{d}^{G^{\prime}}(u)\}_{u\in V}$ as the output. This finishes the algorithm description, which uses 11 call to $\mathcal{A}$} with additional $\widetilde{O}\left(m\right)$ work and $\widetilde{O}\left(1\right)$ depth.
@@ 4.3 Parallel Exact Max Flow via $\tilde{O(}1)$-Approximate Congestion Shortcuts
For a shortcut algorithm $\mathcal{A}$} which outputs a shortcut $G^{\prime}$ of the input graph $G=(V,E)$, a backward mapping algorithm for $\mathcal{A}$} takes input $G$ and $G^{\prime}$ generated by $\mathcal{A}$}, along with a flow $F^{\prime}$ in $G^{\prime}$ with congestion 11, outputs a flow $F$ in $G$ routing the same demand as $F^{\prime}$ with congestion $\kappa$.
@@
The algorithm takes an input graph of $m$ edges and $n$ vertices, uses 11 call to $\mathcal{A}$} and $\mathcal{A}$$^{\prime}$, and additional $\widetilde{O}\left(\mathrm{poly}(h)\cdot(m+m^{\prime})\right)$ work and $\mathrm{poly}(h)$ depth using Lemma 4.12.
@@ C.1 Construction of [VXX24] Lower Bound
  • Edges $E(G)$. Before constructing the edge set of $G$, we define a permutation $\sigma$ of [0,r1][0,r-1] as follows. For any $i\in[0,r-1]$, we define the $i$th element $\sigma_{i}$ of $\sigma$ as the integer whose corresponding $\lg(r)$-bit binary representation can be reversed to obtain the $\lg(r)$-bit binary representation of ii. Let $v_{0},\dots,v_{2r-1}$ be a sequence of vectors where

### v2

@@ 4.2 Exact Shortest Paths via $(1+\epsilon)$-Approximate Distance Shortcuts
We first present an algorithm for computing $(1+o(1/\log n))$-approximate SSSP in an $O(m)$-edge, $O(n)$-vertex directed graph with $1$ call to $\mathcal{A}$} and additional $\widetilde{O}\left(m+m^{\prime}\right)$ time and $\widetilde{O}\left(h\right)$ depth. We then show how to solve exact SSSP of an $m$-edge $n$-vertex graph using $\widetilde{O}\left(1\right)$ calls to the approximation algorithm with additional $\widetilde{O}\left(m\right)$ work and $\widetilde{O}\left(h\right)$ depth. Combining the two reductions finishes the proof.
@@
We can compute an $h$-limited $(1+o(1/\log n))$-approximate shortest path distances $\{\tilde{d}^{G^{\prime}}(u)\}_{u\in V}$ in $\widetilde{O}\left(m+m^{\prime}\right)$ work and $\widetilde{O}\left(h\right)$ depth (by setting the parameter $p=m/k$ in their paper). We use $\{\tilde{d}^{G^{\prime}}(u)\}_{u\in V}$ as the output. This finishes the algorithm description, which uses $1$ call to $\mathcal{A}$} with additional $\widetilde{O}\left(m\right)$ work and $\widetilde{O}\left(1\right)$ depth.
@@ 4.3 Parallel Exact Max Flow via $\tilde{O(}1)$-Approximate Congestion Shortcuts
For a shortcut algorithm $\mathcal{A}$} which outputs a shortcut $G^{\prime}$ of the input graph $G=(V,E)$, a backward mapping algorithm for $\mathcal{A}$} takes input $G$ and $G^{\prime}$ generated by $\mathcal{A}$}, along with a flow $F^{\prime}$ in $G^{\prime}$ with congestion $1$, outputs a flow $F$ in $G$ routing the same demand as $F^{\prime}$ with congestion $\kappa$.
@@
The algorithm takes an input graph of $m$ edges and $n$ vertices, uses $1$ call to $\mathcal{A}$} and $\mathcal{A}$$^{\prime}$, and additional $\widetilde{O}\left(\mathrm{poly}(h)\cdot(m+m^{\prime})\right)$ work and $\mathrm{poly}(h)$ depth using Lemma 4.12.
@@ C.1 Construction of [VXX24] Lower Bound
  • Edges $E(G)$. Before constructing the edge set of $G$, we define a permutation $\sigma$ of [0,r1][0,r-1] as follows. For any $i\in[0,r-1]$, we define the $i$th element $\sigma_{i}$ of $\sigma$ as the integer whose corresponding $\lg(r)$-bit binary representation can be reversed to obtain the $\lg(r)$-bit binary representation of $i$. Let $v_{0},\dots,v_{2r-1}$ be a sequence of vectors where